OpenStudy (anonymous):

Is the following relation a function? y equals the sine of 3 x Yes No

OpenStudy (anonymous):

anybocy minding to help me? @robtobey

OpenStudy (freckles):

do you know what a function is ?

OpenStudy (freckles):

Or do you know any tests the relation must pass to be considered a function?

OpenStudy (anonymous):

A function is an expression involving one or more variables

OpenStudy (freckles):

Are you sure that just that is the definition of a function?

OpenStudy (anonymous):

yes that was the one i was taught

OpenStudy (anonymous):

oh yeah it also relates to an input and output right?

OpenStudy (freckles):

A relation is a function if for every x element of the domain, there is a unique y element of the range assigned to each x. Example: y=x^2 is a function because each x of the domain is assigned one y value of the range y^2=x is not a function. Why? when x=1, y=1 or -1. So x is assigned two different y-values here. We can only have one unique y assigned to each element of the domain.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

I still need help with the question

OpenStudy (freckles):

Yeah that is what I'm trying to do. In order for you to know how to do the question I think you need to know the definition of a function.

OpenStudy (freckles):

Or at least any tests that come with that function.

OpenStudy (freckles):

so it should be clear from the definition of a function that in order for a relation to be considered a function it must pass the vertical line test.

OpenStudy (freckles):

Does your relation pass the vertical line test?

OpenStudy (anonymous):

NO?

OpenStudy (anonymous):

I was not yelling srry did note mean to put the caps on

OpenStudy (freckles):

Just curious...Why is your answer no?

OpenStudy (freckles):

Have you graphed y=sin(3x)?

OpenStudy (anonymous):

on the graph it is not vertical there are just swirly lines?

OpenStudy (freckles):